Mary's age is twice Fred's age and one-third of Sally's age. If the sum of their ages is 99,how old are the three friends ?
Answer by Theo(13342)

let:

m = mary's age
f = fred's age
s = sally's age.

m = 2*f
m = s/3

this leads to:

f = m/2
s = 3*m

sume of their ages is 99.

m + f + s = 99

substitute for f and s to get:

m + m/2 + 3*m = 99

combine like terms to get:

4*m + m/2 = 99

since m/2 = .5*m, this becomes:

4.5 * m = 99

divide 99 by 4.5 to get m = 22

f = m/2 = 11

s = 3*m = 66

their ages are:

m = 22
f = 11
s = 66

sum of their ages is 99.

mary is 22 years old
fred is 11 years old
sally is 66 years old